OpenStudy (anonymous):

A particle is moving along a staright line decelerates uniformly from 40cm/s to 20cm/s and then has a constant acceleration of 20m/s2 during the next 4sec. The average speed during the whole time interval in cm/s is a-57 b-140 c-86 d-43

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

In your question: a) Should "......acceleration of 20m/s2........" have been written as "...acceleration of 20cm/s^2......" ? b) The time taken to decelerate uniformly from 40cm/s to 20cm/s has not been given. Please check your posted question against the original question.

OpenStudy (anonymous):

Time is 5 sec and am sorry it is cm/s^2

OpenStudy (kropot72):

The average speed during the deceleration is given by: \[\frac{40+20}{2}=30\ m/s\] The distance traveled during the deceleration is \[30\times5=150\ cm\] The final speed at the end of the acceleration is \[20+(4\times20)=100\ m/s\] The average speed during the acceleration is \[\frac{20+100}{2}=60\ m/s\] The distance traveled during the acceleeration is \[60\times4=240\ cm\] The total distance traveled is 150 + 240 = 390 cm in a total of 5 + 4 = 9 seconds. Now use the formula \[speed=\frac{distance}{time}\] to find the solution.

OpenStudy (anonymous):

For average speed which formula do you use ? Can ya right it in general form and Thank you So MUCH ^_^

OpenStudy (anonymous):

YAY \O/ I got the answer its D . ^_^ Thanks a bunch :D

OpenStudy (kropot72):

Your choice is correct! Good work :)

OpenStudy (anonymous):

but wait please tell me the general formula for average speed So it can help me

OpenStudy (anonymous):

in future *

OpenStudy (anonymous):

40+20/2=30 m/s total Acceleration/time = average speed ???/

OpenStudy (kropot72):

For motion in a straight line, when the speed changes uniformly with time, the average speed in any time interval equals one-half the sum of the speeds at the beginning and at the end of the interval. Therefore the average speed v-bar between t = 0 and t = t is \[\bar {v}=\frac{v _{0}+v}{2}\]

OpenStudy (anonymous):

oh got it . Thank you so much ya are superb . Such a awesome explanation ^_^

OpenStudy (kropot72):

So , using this formula for the deceleration time interval: \[v _{0}=40\ m/s\] \[v=20\ m/s\] \[\bar {v}=\frac{40+20}{2}=30\ m/s\]

OpenStudy (anonymous):

And isn't it cm/s instead of m/s in average speed ??

OpenStudy (kropot72):

My bad :( Good spotting!

OpenStudy (anonymous):

Thank you so much ^_^

OpenStudy (kropot72):

You're welcome :)